OpenStudy (anonymous):

lim n-->infinity, (3^n + 2^n)/6^n

OpenStudy (anonymous):

U can write it as (1/2)^n + (1/3)^n

OpenStudy (anonymous):

Now, 1^(anything) =1

OpenStudy (anonymous):

and 1^n as n---> infinity = infinity That should help

OpenStudy (anonymous):

Start by rewriting the bottom into 2^n+3^n. Now, you can rewrite the whole expression into a format like:

OpenStudy (anonymous):

\[\frac{ a+b }{ab }\]

OpenStudy (anonymous):

which is the same thing as: \[\frac{ a }{ ab }+\frac{ b }{ab}\]

OpenStudy (anonymous):

so you end up with 1/a + 1/b, where if you take the limit as x-> infinity, it approaches...you finish the rest :)